y+ 2y + y = 0, y(0) = 1 and y(1) = 3 Solve the initial-value differential equation y+ 4y + 4y = 0 subject to the initial


Mathematical Physics 2 H.W.4 J.+y-6y=0 y+ 4y + 4y = 0 y+y=0 Subject to the initial conditions (0) = 2 and y(0) = 1 y-
Answer #1

y+ 2yty=0 we take solution of the form the a emx than : ys meme, y=m2 emix than (m27amt 1)=0 (m+1)=0 m=-1,-1 than two solSecond ? +ylthy zo than. m2+4m +4)=0 Am+a)=0 m=-2-2 y = 72 A+XB) At y(0)=2 Y(0) = A=2. ay= -QA e B Bezze_ge Beze 1. At o.
Answer #2
y"+2y'+y=0